OpenStudy (anonymous):

factor 6x^2-x-1 then tell meh what you did lol.

OpenStudy (anonymous):

6x^2-3x+2x-1 =?

OpenStudy (anonymous):

i factored by grouping

OpenStudy (anonymous):

(2 x-1) (3 x+1) Grouping

OpenStudy (anonymous):

oh totally.

myininaya (myininaya):

ax^2+bx+c a=6 b=-1 c=-1 find factors of a*c that have product a*c and have sum b a*c=6(-1)=-6=-2(3)=2(-3) b=-1=-3+2 bx=-3x+2x replace! 6x^2-3x+2x-1 now factor by grouping 3x(2x-1)+(2x-1) (2x-1)(3x+1)

OpenStudy (anonymous):

^

myininaya (myininaya):

satellite you factored by grouping? i'm prode of you :)